What is the molarity of 49 g of H2SO4 dissolved in 0.3 L of solution?


1
Expert's answer
2021-04-02T05:19:41-0400

Solution.

"m=49g;"

"M(H_2SO_4)=98g\/mol;"

"V=0.3L;"

"M=\dfrac{n}{V};"

"n=\dfrac{m}{M(H_2SO_4)};"

"n=\dfrac{49g}{98g/mol}=0.5mol;"

"M=\dfrac{0.5mol}{0.3L}=1.67M;"


Answer: "M=1.67M;"